What is the formula for the following sequence:
12, 42, 92, 162,...

Found 2 solutions by dabanfield, richwmiller:
Answer by dabanfield(803)

12, 42, 92, 162,...
The first digit in each case is 1^2=1, 2^2=4 3^2=9 and 4^2=16.
The second digit is always 2.
So the next number should be 5^2 followed by 2, that is 252.

Answer by richwmiller(17219)

a_n = 2*(5n^2+1)
Sorry but 16 is not the first digit of 162
2*(5*5^2+1)=2*(5*25+1)
2*(126)=252